OpenStudy (anonymous):

If a regular polygon has exterior angles that measure approximately 17.14º each, how many sides does the polygon have?

OpenStudy (perl):

formula for exterior angle of a regular polygon $$ \Large { \rm exterior~ angle = \frac{360}{n} } $$

OpenStudy (perl):

now plug in the formula and solve for n

OpenStudy (anonymous):

So it would be something like, \[Exterior Angle = 360/17.14\]

OpenStudy (wolf1728):

I'd say do the division and you have your answer.

OpenStudy (wolf1728):

Yes, divide 360 by 17.14 and what do you get?

OpenStudy (perl):

$$ \Large \rm n = \frac{360}{17.14} $$

OpenStudy (wolf1728):

I guess lildunkboi isn't here anymore.

OpenStudy (anonymous):

Sorry Internet went out and didn't put my replies through